An object falls a distance h from rest. If it travels 0.440h in the last 1.00 s, find (a) the time and (b) the height of its fall.
Answer #1

Gravitational acceleration = g = 9.81 m/s2

Initial speed of the object = V0 = 0 m/s (At rest)

Initial height of the object = H

Time taken by the object to reach the ground = T

H = V0T + gT2/2

H = (0)T + gT2/2

H = gT2/2

Speed of the object 1 sec before hitting the ground = V1

V1 = V0 + g(T - 1)

V1 = 0 + g(T - 1)

V1 = gT - g

Speed of the object just before hitting the ground = V2

V2 = V0 + gT

V2 = 0 + gT

V2 = gT

Distance through which the object falls in the last second = L = 0.44H = 0.44(gT2/2) = 0.22gT2

V22 = V12 + 2gL

(gT)2 = (gT - g)2 + 2g(0.22gT2)

g2T2 = g2T2 - 2g2T + g2 + 0.44g2T2

T2 = T2 - 2T + 1 + 0.44T2

0.44T2 - 2T + 1 = 0

T = 3.97 sec or 0.57 sec

Time period cannot be lower than 1 sec.

T = 3.97 sec

H = gT2/2

H = (9.81)(3.97)2/2

H = 77.3 m

a) Time of the fall = 3.97 sec

b) Height of the fall = 77.3 m